Ticket Number: HEA-NOC/19980512-01 Ticket Status: CLOSE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEAnet <==> TCD
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 19980512 17:21 UTC+1
Ticket Closed: 19981028 17:04:27 UTC Problem End: 19980512 17:37 UTC+1
The line from HEAnet to Trinity College, Dublin was unreliable,
apparently due to a clocking issue with the Micromuxes at each end.
Performance to and from TCD was impeded.
Reseated the BNC connectors at the HEAnet end; this reset the line
and cleared the problem.
Time to Fix: